Wayne is 20 years older than Sabrina. In 8 years, Wayne will be twice as old as Sabrina. How old is each now?
Answer by macston(5194)

W=Wayne; S=Sabrina=W-20yrs;
W+8yrs=2(S+8yrs) Substitute for S.
W+8yrs=2(W-20yrs+8yrs)
W+8yrs=2W-24yrs
32 yrs=W ANSWER 1: Wayne is 32 years old.
S=W-20yrs=32yrs-20yrs=12 yrs ANSWER 2: Sabrina is 12 years old.
CHECK:
W+8yrs=2(S+8yrs)
32yrs+8yrs=2(12yrs+8yrs)
40yrs=40yrs